Consider A, B and C to be an events. What is the sense of the following equations:
a) ABC = A;
b) A+B+C = A?
1
Expert's answer
2010-07-06T10:49:13-0400

a) ABC = A - The simultaneous occurrence of the events A, B and C is equivalent to the event A.
b) A+B+C = A - The occurrence of at least one of the events A, B or C is equivalent to the event A.